OpenStudy (anonymous):

x=2(mod 3) x=3(mod 7) x=7(mod 5) find a unique incongruent solution to modulo n. (number theory problem using chinese remainder theorem) please help!!

OpenStudy (anonymous):

Yes this is indeed a basic application of CRT, you want me to do it for you?

OpenStudy (anonymous):

that would be nice11

OpenStudy (anonymous):

\[ x=17 \]

OpenStudy (anonymous):

i don't think so.. because based on the book, the answer would be in the pattern a=b(mod n)

OpenStudy (anonymous):

Use your head man, this is elementary.

OpenStudy (mathmate):

First, check if there was a typo in the question, x=7(mod5) doesn't not sound right. Will proceed as if it were x=2(mod5). Method does not change. The basic requirement is to satisfy all conditions x=2(mod 3) x=3(mod 7) x=2(mod 5) We realize that there are infinitely many solution, all of which satisfy all three conditions. Suppose 17 is one such solution [check: 17=2(mod3), 17=3(mod7), 17=2(mod5), OK]. We calculate the LCM of 3,7 and 5 to be 3*7*5=105. In other words, 105 has a remainder of 0 when divided by 3, 7 or 5. Hence 17 mod 105 is a general solution to the above problem, since if we add multiples of 105 to 17, the remainder will not change.

OpenStudy (anonymous):

"First, check if there was a typo in the question, x=7(mod5) doesn't not sound right. Will proceed as if it were x=2(mod5). " --- Why? As far as I can see (3,4,5)=1 so perfect candidate for CRT.

OpenStudy (mathmate):

I suspected a typo because we don't usually write x=7(mod5) where 7>5.

OpenStudy (anonymous):

That's true :)

OpenStudy (mathmate):

:)